OpenStudy (kendricklamar2014):

A bag contains 7 green marbles and 4 white marbles. You select a marble at random. What are the odds in favor of picking a green marble? A. 7:11 B. 7:4 C. 4:7 D. 3:7 Thinking A or B?

OpenStudy (anonymous):

Is there are a total of 11 marbles in a bag and there are 7 green, your odds would be... Remember that a/b can also be written as a:b.

OpenStudy (kendricklamar2014):

Thanks

OpenStudy (kropot72):

The odds are the probability of an event occurring to that of its not occurring: P(green)=7/11 P(white) = 4/11 The required odds are therefore 7:4

OpenStudy (kendricklamar2014):

Is it A or B?

OpenStudy (kendricklamar2014):

@YanaSidlinskiy

OpenStudy (kropot72):

The question asks for the odds of a green marble. Choice B is the odds. Choice A gives the probability of green, which is not the same thing as the odds in favor of green.

OpenStudy (kendricklamar2014):

thanks, that what i was thinking :)

OpenStudy (kropot72):

You're welcome :)